### v1

@@ 3.1 SHAP Variance Decomposition
    SHAP Linearity: $\phi_{i}$ is linear in tree outputs (holds for TreeSHAP)
Law of Total Variance
Apply the law of total variance (first decomposition) [60] conditioned on $f$:
$\text{Var}_{D,f}(\phi_{i})=\mathbb{E}_{f}[\text{Var}_{D}(\phi_{i}|f)]+\text{Var}_{f}(\mathbb{E}_{D}[\phi_{i}|f])$ (2)
This gives the standard aleatoric (first term) and epistemic (second term) decomposition, but ignores the model-data dependency.
Refinement for Entanglement
The term $\mathbb{E}_{f}[\text{Var}_{D}(\phi_{i}|f)]$ is decomposed by conditioning on $D$:
$\begin{split}\mathbb{E}_{f}[\mathrm{Var}_{D}(\phi_{i}|f)]&=\mathbb{E}_{D}[\mathrm{Var}_{f}(\phi_{i}|D)]\\ &\quad+\left(\mathbb{E}_{f}[\mathrm{Var}_{D}(\phi_{i}|f)]-\mathbb{E}_{D}[\mathrm{Var}_{f}(\phi_{i}|D)]\right)\end{split}$ (3)
The excess term arises from non-commutativity of expectations due to $P(f,D)\neq P(f)P(D)$.
Covariance Identification
The entanglement term emerges from:
$\mathcal{C}(f,D)=\text{Cov}\left(\mathbb{E}_{D}[\phi_{i}|f],\text{Var}_{f}(\phi_{i}|D)\right)$ (4)
Derivation:
@@
    Expand $\mathbb{E}_{f}[\text{Var}_{D}(\phi_{i}|f)]$ using iterated expectation:
    $\begin{split}\mathbb{E}_{f}[\text{Var}_{D}(\phi_{i}|f)]=\mathbb{E}_{D}[\text{Var}_{f}(\phi_{i}|D)]+\\ \text{Cov}\left(\mathbb{E}_{D}[\phi_{i}|f],\text{Var}_{f}(\phi_{i}|D)\right)\end{split}$ (5)
@@
    Substitute into total variance:
    $\text{Var}_{D,f}(\phi_{i})=\mathbb{E}_{D}[\text{Var}_{f}(\phi_{i}|D)]+\text{Var}_{f}(\mathbb{E}_{D}[\phi_{i}|f])+2\mathcal{C}(f,D)$ (6)
Special Case: Tree Ensembles
For Random Forests with $B$ trees trained on bootstrap samples $\{D_{b}\}$:
  1. 1.
    Aleatoric Term:
    $\mathbb{E}_{D}[\text{Var}_{f}(\phi_{i}|D)]\approx\frac{1}{B}\sum_{b=1}^{B}\text{Var}_{T\in\mathcal{T}_{b}}(\phi_{i}^{(T)})$ (7)
    where $\mathcal{T}_{b}$ are trees trained on $D_{b}$.
  2. 2.
    Epistemic Term:
    $\text{Var}_{f}(\mathbb{E}_{D}[\phi_{i}|f])\approx\text{Var}_{b}\left(\frac{1}{|\mathcal{T}_{b}|}\sum_{T\in\mathcal{T}_{b}}\phi_{i}^{(T)}\right)$ (8)
@@
    Entanglement Term:
    $\mathcal{C}(f,D)\propto\sum_{b=1}^{B}\left(\bar{\phi}_{i}^{(b)}-\bar{\phi}_{i}\right)\left(\sigma_{i}^{(b)2}-\bar{\sigma}_{i}^{2}\right)$ (9)
@@
    • $\bar{\phi}_{i}^{(b)}=$ mean SHAP for trees in bootstrap $b$
    • $\sigma_{i}^{(b)2}=$ SHAP variance for trees in bootstrap $b$
The UbiQTree estimator approximates this decomposition via:
@@
    E-SHAP’s Dirichlet-weighted sampling preserves this covariance structure, unlike bootstrap methods that assume $P(f,D)\approx P(f)P(D)$
@@
3.2 Evidence Theory: Tree Ensembles
The Dempster-Shafer theory (DST) is a mathematical framework for reasoning under uncertainty, particularly when evidence is incomplete, imprecise, or conflicting. DST assigns belief masses to sets or intervals of possible outcomes, allowing for the explicit representation of ignorance and epistemic uncertainty. Key DST concepts include belief mass ($m$), belief ($Bel$), plausibility ($Pl$), and ignorance. DST is widely used in artificial intelligence, sensor fusion, medical diagnostics, risk assessment, and autonomous systems, where managing uncertainty and combining evidence from multiple sources is critical. It provides a systematic and flexible approach to uncertainty, enabling AI and decision systems to model ignorance and combine evidence in ways that classical probability theory cannot. DST is a valuable tool for managing uncertainty and combining evidence in AI and decision systems. [61, 53].
@@ 3.2 Evidence Theory: Tree Ensembles
For a tree ensemble with $K$ trees, the Basic Probability Assignment (BPA) for SHAP value $\phi_{i}$ belonging to interval $A\subseteq\mathbb{R}$ is:
$m(A)=\frac{1}{K}\sum_{k=1}^{K}\mathbb{I}\left(\phi_{i}^{(k)}\in A\right)$ (10)
where $\phi_{i}^{(k)}$ is the SHAP value from tree $T_{k}$. The Belief and Plausibility functions satisfy:
@@
Proof: BPA Construction Each tree represents an independent evidence source. The BPA is the proportion of trees supporting interval $A$, satisfying:
@@
  • $\sum_{A\subseteq\mathbb{R}}m(A)=1$ (normalization)
@@
For nested intervals $A_{1}\subseteq A_{2}\subseteq\cdots\subseteq A_{n}$:
$\text{Bel}(A_{n})=\sum_{j=1}^{n}m(A_{j})\quad\text{(consonant structure)}$ (12)
This follows from the definition of Belief as the total evidence supporting $A$.
@@
For conflicting explanations (e.g., positive vs. negative impact):
$\text{Pl}(A)-\text{Bel}(A)=1-\sum_{B\subseteq A}m(B)-\sum_{B\subseteq A^{c}}m(B)$ (13)
where $A^{c}$ is the complement. This quantifies the probability mass assigned to sets overlapping both $A$ and $A^{c}$.
@@
Since trees are exchangeable:
@@
By the Law of Large Numbers, Belief converges to the true probability.
@@
$\mathcal{C}_{i}=\sup_{A\subseteq\mathbb{R}}\left[\text{Pl}(A)-\text{Bel}(A)\right]$ (15)
which measures the maximum ambiguity in SHAP assignments.
3.3 Uncertainty Theory: Application to SHAP
Uncertainty theory by Liu et al. [62, 63] is a mathematical framework designed to address epistemic uncertainty arising from incomplete knowledge, small sample sizes, or reliance on expert judgment. The theory is based on four axioms: normality, monotonicity, self-duality, and countable subadditivity. The central concept is the uncertainty distribution, denoted by the symbol $\Gamma\colon\mathbb{R}\to[0,1]$, which quantifies the degree of belief that a variable takes on values less than or equal to $\phi_{i}$. This distribution is characterized by a value of 0 for implausible values and a value of 1 for fully plausible values. It is also monotonically increasing as values become more plausible. Uncertainty distributions model subjective confidence rather than frequency or likelihood. This makes them useful in situations with limited or non-statistical data. When applied to SHAP or feature attribution in AI, an uncertainty distribution can represent confidence in a feature’s attribution magnitude. This allows practitioners to explicitly model and quantify their uncertainty about the importance of each feature, especially when data is scarce or unreliable. Entropy minimization can guide optimal data acquisition, thereby improving model interpretability and reliability [64, 62, 63, 65, 66].
Theorem 2 (Uncertainty Distribution).
@@ 3.3 Uncertainty Theory: Application to SHAP
Lemma 1 (Optimal Acquisition).
@@
3.4 Dirichlet Process Hypothesis Sampling
Dirichlet processes (DPs) are key to Bayesian nonparametric modeling. They allow for flexible inference over distributions with unknown and potentially infinite underlying clusters. DPs are useful for modeling uncertainty in complex spaces like SHAP values and their clusters across tree ensembles. Each sample from a DP is a discrete probability distribution. DPs are parameterized by a concentration parameter, $\alpha$, and a base distribution, $G_{0}$. In mixture models, DPs allow for an unbounded number of mixture components, thereby adapting model complexity to the data. In tree ensembles, each tree is a hypothesis about feature attributions. By modeling the distribution of SHAP values across trees with a DP mixture model, one can cluster SHAP values without specifying the number of clusters or modes in advance. This method captures both diversity and epistemic uncertainty in feature attributions due to model variability and quantifies uncertainty by examining the posterior distribution of clusters or modes of SHAP values. This provides richer uncertainty estimates than standard bootstrap or ensemble variance methods. DPs have several advantages over parametric and bootstrap methods. First, they avoid the need to fix the number of clusters or modes in advance. DPs adapt to model complexity as more data or trees are considered. DPs mitigate under- or overfitting and provide a more nuanced, probabilistic view of uncertainty in SHAP attributions. DPs are widely used in machine learning for clustering and mixture models where the number of components is unknown. Such an approach allows for a richer and more flexible quantification of epistemic uncertainty in SHAP attributions by leveraging the full power of Bayesian nonparametrics [67, 68, 69, 70, 71, 72].
Theorem 3 (Constructing the Dirichlet Process).
@@ 3.4 Dirichlet Process Hypothesis Sampling
    By the de Finetti theorem, infinite exchangeable trees induce a Gaussian process [73, 74, 75, 76].
@@
    The Dirichlet process is the de Finetti measure for Pólya sequences [77].
@@
    SHAP values are continuous linear operators, preserving convergence [17].
4 Methodology

### v2

@@ INTRODUCTION
    SHAP Additivity: $\phi_{i}$ is additive in tree outputs (holds for TreeSHAP in Random Forests)
  4. 4.
    Regular Conditional Distributions: $\mathbb{E}_{D}[\phi_{i}|f]$ and
    $\mathbb{E}_{f}[\phi_{i}|D]$ exist as proper random variables
Step 1: Total Variance Definition
Apply the law of total variance (first decomposition) 9
$\text{Var}_{D,f}(\phi_{i})$ $=\mathbb{E}_{D,f}[\phi_{i}^{2}]-\left(\mathbb{E}_{D,f}[\phi_{i}]\right)^{2}$
Step 2: Expand $\mathbb{E}_{D,f}[\phi_{i}^{2}]$ using Law of Total Expectation
$\mathbb{E}_{D,f}[\phi_{i}^{2}]$ $=\mathbb{E}_{D}\left[\mathbb{E}_{f}[\phi_{i}^{2}|D]\right]$
$=\mathbb{E}_{D}\left[\text{Var}_{f}(\phi_{i}|D)+\left(\mathbb{E}_{f}[\phi_{i}|D]\right)^{2}\right]$
$=\mathbb{E}_{D}[\text{Var}_{f}(\phi_{i}|D)]+\mathbb{E}_{D}\left[(\mathbb{E}_{f}[\phi_{i}|D])^{2}\right]$
Step 3: Expand $\left(\mathbb{E}_{D,f}[\phi_{i}]\right)^{2}$
$\left(\mathbb{E}_{D,f}[\phi_{i}]\right)^{2}$ $=\left(\mathbb{E}_{f}\left[\mathbb{E}_{D}[\phi_{i}|f]\right]\right)^{2}$
$=\mathbb{E}_{f}\left[(\mathbb{E}_{D}[\phi_{i}|f])^{2}\right]-\text{Var}_{f}(\mathbb{E}_{D}[\phi_{i}|f])$
This follows from the identity: $\left(\mathbb{E}[X]\right)^{2}=\mathbb{E}[X^{2}]-\text{Var}(X)$.
Step 4: Combine Terms
$\text{Var}_{D,f}(\phi_{i})$ $=\mathbb{E}_{D}[\text{Var}_{f}(\phi_{i}|D)]+\mathbb{E}_{D}\left[(\mathbb{E}_{f}[\phi_{i}|D])^{2}\right]$
$\quad-\mathbb{E}_{f}\left[(\mathbb{E}_{D}[\phi_{i}|f])^{2}\right]+\text{Var}_{f}(\mathbb{E}_{D}[\phi_{i}|f])$
Step 5: Identify the Entanglement Term:
Let us examine the difference:
$\Delta$ $=\mathbb{E}_{D}\left[(\mathbb{E}_{f}[\phi_{i}|D])^{2}\right]-\mathbb{E}_{f}\left[(\mathbb{E}_{D}[\phi_{i}|f])^{2}\right]$
Define the covariance between model-centric and data-centric expectations:
$\mathcal{C}(f,D)$ $=\text{Cov}\left(\mathbb{E}_{D}[\phi_{i}|f],\mathbb{E}_{f}[\phi_{i}|D]\right)$
$=\mathbb{E}_{f,D}\left[\mathbb{E}_{D}[\phi_{i}|f]\cdot\mathbb{E}_{f}[\phi_{i}|D]\right]$
$\quad-\mathbb{E}_{f,D}[\mathbb{E}_{D}[\phi_{i}|f]]\cdot\mathbb{E}_{f,D}[\mathbb{E}_{f}[\phi_{i}|D]]$
Observe that:
$\mathbb{E}_{f,D}\left[\mathbb{E}_{D}[\phi_{i}|f]\cdot\mathbb{E}_{f}[\phi_{i}|D]\right]$ $=\mathbb{E}_{D}\left[(\mathbb{E}_{f}[\phi_{i}|D])^{2}\right]$
$\mathbb{E}_{f,D}\left[(\mathbb{E}_{D}[\phi_{i}|f])^{2}\right]$ $=\mathbb{E}_{f}\left[(\mathbb{E}_{D}[\phi_{i}|f])^{2}\right]$
Therefore:
$\Delta$ $=\mathbb{E}_{D}\left[(\mathbb{E}_{f}[\phi_{i}|D])^{2}\right]-\mathbb{E}_{f}\left[(\mathbb{E}_{D}[\phi_{i}|f])^{2}\right]$
$=-\mathcal{C}(f,D)$
Step 6: Final Decomposition:
Substituting back the values, we have:
$\text{Var}_{D,f}(\phi_{i})$ $=\mathbb{E}_{D}[\text{Var}_{f}(\phi_{i}|D)]+\text{Var}_{f}(\mathbb{E}_{D}[\phi_{i}|f])-\mathcal{C}(f,D)$
Thus we have the decomposition:
$\boxed{\text{Var}_{D,f}(\phi_{i})=\mathbb{E}_{D}[\text{Var}_{f}(\phi_{i}|D)]+\text{Var}_{f}(\mathbb{E}_{D}[\phi_{i}|f])-\mathcal{C}(f,D)}$ (2)
where
$\mathcal{C}(f,D)=\text{Cov}\left(\mathbb{E}_{D}[\phi_{i}|f],\mathbb{E}_{f}[\phi_{i}|D]\right)$ (3)
Interpretation of the Entanglement Term
The entanglement term $\mathcal{C}(f,D)$ measures the covariance between:
  • Model-centric expectation: $\mathbb{E}_{D}[\phi_{i}|f]$ – average SHAP across datasets for a fixed model
  • Data-centric expectation: $\mathbb{E}_{f}[\phi_{i}|D]$ – average SHAP across models for a fixed dataset
Key Insights:
@@
    Negative Sign: When model and data "agree" on feature importance (positive covariance), the total variance decreases because the uncertainty sources are aligned.
@@
    Non-commutativity: The term arises precisely because $\mathbb{E}_{D}[\mathbb{E}_{f}[\cdot|D]]\neq\mathbb{E}_{f}[\mathbb{E}_{D}[\cdot|f]]$ when $P(f,D)\neq P(f)P(D)$.
@@
    Practical Meaning: In tree ensembles, this occurs when certain data distributions consistently produce models that assign similar feature importance patterns.
Special Case: Independent Model and Data
If $P(f,D)=P(f)P(D)$ (model training independent of data sampling), then:
$\mathbb{E}_{D}[\mathbb{E}_{f}[\phi_{i}|D]]=\mathbb{E}_{f}[\mathbb{E}_{D}[\phi_{i}|f]]=\mathbb{E}_{D,f}[\phi_{i}]$
and the covariance term vanishes:
$\mathcal{C}(f,D)=0$
reducing to the standard decomposition:
$\text{Var}_{D,f}(\phi_{i})=\mathbb{E}_{D}[\text{Var}_{f}(\phi_{i}|D)]+\text{Var}_{f}(\mathbb{E}_{D}[\phi_{i}|f])$
Empirical Estimation for Tree Ensembles
For Random Forests with $B$ bootstrap samples, the entanglement term can be estimated as:
$\hat{\mathcal{C}}(f,D)=\frac{1}{B}\sum_{b=1}^{B}\left(\bar{\phi}_{i}^{(b)}-\bar{\phi}_{i}\right)\left(\bar{\phi}_{i}^{(b)^{\prime}}-\bar{\phi}_{i}\right)$
@@
  • $\bar{\phi}_{i}^{(b)}=\frac{1}{|\mathcal{T}_{b}|}\sum_{T\in\mathcal{T}_{b}}\phi_{i}^{(T)}$: mean SHAP for trees in bootstrap $b$ (data-centric)
  • $\bar{\phi}_{i}^{(b)^{\prime}}=\frac{1}{B}\sum_{j=1}^{B}\phi_{i}^{(T_{j})}$ where $T_{j}$ are trees from method $b$ across bootstraps (model-centric)
  • $\bar{\phi}_{i}=\frac{1}{B}\sum_{b=1}^{B}\bar{\phi}_{i}^{(b)}$: overall mean SHAP
Theoretical Implications
This decomposition reveals that the standard uncertainty quantification framework overestimates total variance when model and data expectations are positively correlated. The entanglement term acts as a correction factor that accounts for the dependency structure between model training and data sampling processes.
For high-stakes AI systems, this refined understanding enables more precise uncertainty attribution and better reliability assessment of feature importance explanations. The UbiQTree estimator approximates this decomposition via:
@@
    Dirichlet-weighted sampling preserves this covariance structure, unlike bootstrap methods that assume $P(f,D)\approx P(f)P(D)$
@@
Evidence Theory: Tree Ensembles
The Dempster-Shafer theory (DST) is a mathematical framework for reasoning under uncertainty, particularly when evidence is incomplete, imprecise, or conflicting. DST assigns belief masses to sets or intervals of possible outcomes, allowing for the explicit representation of ignorance and epistemic uncertainty. Key DST concepts include belief mass ($m$), belief ($Bel$), plausibility ($Pl$), and ignorance. DST is widely used in artificial intelligence, sensor fusion, medical diagnostics, risk assessment, and autonomous systems, where managing uncertainty and combining evidence from multiple sources is critical. It provides a systematic and flexible approach to uncertainty, enabling AI and decision systems to model ignorance and combine evidence in ways that classical probability theory cannot. DST is a valuable tool for managing uncertainty and combining evidence in AI and decision systems. 78, 77.
@@
For a tree ensemble with $K$ trees, we define a finite frame of discernment by partitioning the range of SHAP values into $M$ disjoint intervals:
$\Omega=\{I_{1},I_{2},\dots,I_{M}\}$
where $I_{j}$ are intervals covering the observed SHAP values with $\bigcup_{j=1}^{M}I_{j}\supseteq[\min(\Phi_{i}),\max(\Phi_{i})]$ and $I_{j}\cap I_{k}=\emptyset$ for $j\neq k$. The Basic Probability Assignment (BPA) for any subset $A\subseteq\Omega$ is:
$m(A)=\frac{1}{K}\sum_{k=1}^{K}\mathbb{I}\left(\phi_{i}^{(k)}\in\textstyle\bigcup_{I\in A}I\right)$ (4)
where $\phi_{i}^{(k)}$ is the SHAP value from tree $T_{k}$ and $\mathbb{I}$ is the indicator function. The Belief and Plausibility functions satisfy:
@@
Proof: BPA Construction Each tree represents an independent evidence source. The BPA is the proportion of trees supporting the union of intervals in $A$, satisfying:
@@
  • $\sum_{A\subseteq\Omega}m(A)=1$ (normalization)
The finite frame ensures all sums are well-defined and finite.
@@
For nested intervals $A_{1}\subseteq A_{2}\subseteq\cdots\subseteq A_{n}$ defined over the real line, we map to our finite frame by taking $A_{j,\Omega}=\{I\in\Omega:I\subseteq A_{j}\}$. Then:
$\text{Bel}(A_{n})=\sum_{B\subseteq A_{n,\Omega}}m(B)\quad\text{(consonant structure)}$ (6)
This follows from the definition of Belief as the total evidence supporting all subsets of $A_{n,\Omega}$.
@@
For conflicting explanations (e.g., positive vs. negative impact), let $A$ be an interval and $A^{c}$ its complement. The ambiguity is:
$\text{Pl}(A)-\text{Bel}(A)=1-\sum_{B\subseteq A_{\Omega}}m(B)-\sum_{B\subseteq A^{c}_{\Omega}}m(B)$ (7)
where $A_{\Omega}=\{I\in\Omega:I\subseteq A\}$ and $A^{c}_{\Omega}=\{I\in\Omega:I\subseteq A^{c}\}$. This quantifies the probability mass assigned to sets overlapping both $A$ and $A^{c}$.
@@
Since trees are exchangeable and the frame is finite:
@@
By the Law of Large Numbers, Belief converges to the true probability when the partition is sufficiently refined.
@@
$\mathcal{C}_{i}=\max_{j=1,\dots,M}\left[\text{Pl}(I_{j})-\text{Bel}(I_{j})\right]$ (9)
which measures the maximum ambiguity in SHAP assignments across the partition intervals.
Theoretical Implications
This finite-frame Dempster-Shafer representation enables rigorous uncertainty quantification for SHAP values in tree ensembles:
  1. 1.
    The finite frame ensures all belief and plausibility sums are finite and well-defined.
  2. 2.
    The partition-based approach aligns with histogram estimation of SHAP distributions.
  3. 3.
    Belief represents confirmed evidence while plausibility represents possible evidence, enabling precise uncertainty decomposition.
  4. 4.
    The conflict measure $\mathcal{C}_{i}$ identifies features with contradictory importance patterns across trees.
For high-stakes AI systems, this evidence-theoretic framework provides a mathematically sound foundation for reliability assessment of feature importance explanations, complementing the variance-based decomposition in Equation (1).
Uncertainty Theory: Application to SHAP
Uncertainty theory by Liu et al. 52, 53 is a mathematical framework designed to address epistemic uncertainty arising from incomplete knowledge, small sample sizes, or reliance on expert judgment. The theory is based on four axioms: normality, monotonicity, self-duality, and countable subadditivity. The central concept is the uncertainty distribution, denoted by the symbol $\Gamma\colon\mathbb{R}\to[0,1]$, which quantifies the degree of belief that a variable takes on values less than or equal to $\phi_{i}$. This distribution is characterized by a value of 0 for implausible values and a value of 1 for fully plausible values. It is also monotonically increasing as values become more plausible. Uncertainty distributions model subjective confidence rather than frequency or likelihood. This makes them useful in situations with limited or non-statistical data. When applied to SHAP or feature attribution in AI, an uncertainty distribution can represent confidence in a feature’s attribution magnitude. This allows practitioners to explicitly model and quantify their uncertainty about the importance of each feature, especially when data is scarce or unreliable. Entropy minimization can guide optimal data acquisition, thereby improving model interpretability and reliability 51, 52, 53, 93, 64.
Theorem 2 (Uncertainty Distribution)
@@
Lemma 1 (Optimal Acquisition)
@@
Dirichlet Process Hypothesis Sampling
Dirichlet processes (DPs) are key to Bayesian nonparametric modeling. They allow for flexible inference over distributions with unknown and potentially infinite underlying clusters. DPs are useful for modeling uncertainty in complex spaces like SHAP values and their clusters across tree ensembles. Each sample from a DP is a discrete probability distribution. DPs are parameterized by a concentration parameter, $\alpha$, and a base distribution, $G_{0}$. In mixture models, DPs allow for an unbounded number of mixture components, thereby adapting model complexity to the data. In tree ensembles, each tree is a hypothesis about feature attributions. By modeling the distribution of SHAP values across trees with a DP mixture model, one can cluster SHAP values without specifying the number of clusters or modes in advance. This method captures both diversity and epistemic uncertainty in feature attributions due to model variability and quantifies uncertainty by examining the posterior distribution of clusters or modes of SHAP values. This provides richer uncertainty estimates than standard bootstrap or ensemble variance methods. DPs have several advantages over parametric and bootstrap methods. First, they avoid the need to fix the number of clusters or modes in advance. DPs adapt to model complexity as more data or trees are considered. DPs mitigate under- or overfitting and provide a more nuanced, probabilistic view of uncertainty in SHAP attributions. DPs are widely used in machine learning for clustering and mixture models where the number of components is unknown. Such an approach allows for a richer and more flexible quantification of epistemic uncertainty in SHAP attributions by leveraging the full power of Bayesian nonparametrics 49, 50, 71, 86, 83, 66.
Theorem 3 (Constructing the Dirichlet Process)
@@
    By the de Finetti theorem, infinite exchangeable trees induce a Gaussian process 29, 31, 30, 73.
@@
    The Dirichlet process is the de Finetti measure for Pólya sequences 60.
@@
    SHAP values are continuous linear operators, preserving convergence 56.
METHODS